OpenStudy (anonymous):

Identify the range of the equation y = −x2 − 12x + 3. All real numbers y ≤ 39 y ≥ −6 y ≤ −105

jimthompson5910 (jim_thompson5910):

\[\large y=-x^2-12x+3\] \[\large y=-(x^2+12x-3)\] \[\large y=-(x^2+12x+36-36-3)\] \[\large y=-((x^2+12x+36)-36-3)\] \[\large y=-((x+6)^2-36-3)\] \[\large y=-((x+6)^2-39)\] \[\large y=-(x+6)^2+39\] Graph is a parabola that opens downward, so the max value of y is 39 So the range is \[y \le 39\] so the answer is choice B